Find the points on the graph of the function f(x)=x+1/x-1. Where the slope of the tangent line is equal to -1/3 determine the equation of the tangent and normal to line
1
Expert's answer
2020-07-07T21:12:26-0400

F(x)=(x+1)/(x-1)

a formula of the derivative f(x)=1*(x-1)-(x+1)*1/(x-1)2=-2/(x-1)2

the derivative of the function is equal to the tangent of the angle of inclination

f(x0)=-1/3=-2/(x0-1)2,

1/6=(x0-1)2,

x01-1=√6 , x01=1+√6,

x02-1=-√6, x02=1-√6,

F(x01)=(2+√6)/(√6) =1+√2/√3,

F(x02)=(2-√6)/(-√6)=1-√2/√3.

Equations of the tangent y=f(x0)(x-x0)+F(x0),

y1=-1/3(x-1-√6)+1+√2/√3,

y2=-1/3(x-1+√6)+1-√2/√3.

Equations of the normal y=-1/f(x0)(x-x0)+F(x0),

y1=3(x-1-√6)+1+√2/√3,

y2=3(x-1+√6)+1-√2/√3.